2. Find the length and width of a rectangular lot with a perimeter of 70 meters if the length is 5 meters more than the width.
Answer by checkley77(12844)

L=W+5
2L+2W=70
2(W+5)+2W=70
2W+10+2W=70
4W=70-10
4W=60
W=60/4
W=15 ANS. FOR THE WIDTH.
L=15+5=20 ANS. FOR THE LENGTH.
pROOF:
2*20+2*15=70
40+30=70
70=70